# Ion Coulomb crystals: an exotic form of condensed matter

### v1

@@ I Introduction
In a seminal article from 1934, Eugene Wigner predicted that conductance electrons in a metal at low densities can crystallize [331]. This is a peculiar situation in condensed matter, where otherwise crystallization is expected at high densities, and it results from the decay of the Coulomb interactions as the inverse of the interparticle distance $a$. For this reason, in an unscreened environment the Coulomb interactions scale with the charge density $n\sim a^{-3}$ as $E_{\rm pot}\sim n^{1/3}$. By comparison, the typical scale of the kinetic energy in a degenerate free electron gas is $E_{\rm kin}\sim\frac{h^{2}}{2ma^{2}}\sim n^{2/3}$ [175]. As a result, at low densities, the Coulomb interactions are dominant over the kinetic energy, leading to crystallization. More generally, at low densities the Wigner crystal phase is favored regardless of the particle statistics, since the overlap of single-particle wave functions is small.
@@
Wigner crystallization, here broadly defined as the formation of ordered arrays of charged particles, can occur in a variety of Coulomb systems. It is expected to occur in the interior of white dwarfs and in the outer crust of neutron stars [323]. It has been experimentally observed with electrons trapped on the surface of liquid helium [128], and in low-dimensional solid state devices including GaAs/GaAlAs heterojunctions [5] and carbon nanotubes [292]. Ordered arrays of charged, solid particles have been experimentally studied with dusty plasmas [312] and colloidal suspensions of polystyrene spheres in water [324]. Wigner crystallization is being pursued through the formation of ultracold neutral plasmas [187, 19] of atomic ions immersed in a background of photo-ionized electrons. In the last three examples, the charged objects interact through a screened Coulomb potential.
Wigner crystallization has also been observed in the laboratory through laser cooling of atomic and molecular ions stored in radiofrequency (rf) Paul traps [78, 332, 84], in Penning ion traps [123], and in ion storage rings [22]. These Wigner crystals of trapped atomic ions have been acquiring increasing importance for quantum technological applications [42, 218] thanks to advances in cooling and trapping [106, 160, 286]. From the point of view of many-body physics, Wigner crystals of trapped ions are unusual examples of solid phases. Crystallization is the result of the interplay between the Coulomb repulsion and the trap potential, and their intrinsic properties are significantly influenced by the confining environment. The ions interact via the long-range Coulomb repulsion, which is not screened by the trap potential, giving rise to dynamical and thermodynamical properties different from the ones of typical crystals in condensed matter.
@@
Wigner crystals of trapped ions, often dubbed ion Coulomb crystals, have been a formidable platform for the simulation of quantum many-body physics [286, 218] and for the realization of milestones of quantum information processing [42]. These advances rely on the detailed knowledge and experimental control of the microscopic dynamics of these strongly-correlated phases of matter [160, 286]. This review reports the progress on the characterization of the equilibrium and out-of-equilibrium dynamics of Wigner crystals of trapped ions, which has enabled this progress and has, in turn, unveiled the properties of an exotic form of condensed matter.
@@
Our starting point is the knowledge summarized in the review by Dubin and O’Neil [98] on the plasma phase and its transition to crystallization. We set our focus on the crystallized phase and discuss the insights gained over two decades of theoretical and experimental investigations. The content here discussed sets the basis for and is complementary to Refs. [160, 286], which report on studies and concepts for exploiting the structure and vibrational spectrum for quantum simulators, and to recent overview articles [83, 313, 213].
@@ I.1 Historical framework and conditions for crystallization
In 1977 Malmberg and O’Neil explained how collections of a single species of trapped charges (ions or electrons) are ideal one-component plasmas  [200]. One-component plasmas (OCP) have been the subject of detailed theoretical investigations for understanding correlations in a diverse range of physical systems [41, 153, 154]. They consist of a single species of point charges that interact through Coulomb repulsion and are immersed in a uniform neutralizing background of given density $n_{0}$. The uniform neutralizing background is assumed to be a continuous, non-polarizable background charge density. In equilibrium, the point charges match their density to that of the neutralizing background. Malmberg and O’Neil showed that the forces that confine ions in a Penning ion trap are mathematically equivalent to a uniform, neutralizing background. More specifically, in a frame rotating with the collection of ions (or electrons) confined in a Penning trap (see Sec. I.2.2), the static thermodynamic properties of the trapped ions are the same as those for an OCP. The argument can be straightforwardly extended to ions confined in radio frequency traps in the so-called pseudo-potential approximation (see Sec. I.2.1).
@@
The coupling constant $\Gamma$ is a measure of the ratio of the Coulomb potential energy between nearest neighbor ions (or more generally point charges) to the average kinetic energy of an ion. Here, $\epsilon_{0}$ is the permittivity of the vacuum, $Q$ is the charge and $T$ is the temperature of the ion, $k_{\mathrm{B}}$ is Boltzmann’s constant, and $a_{\mathrm{WS}}$ is the Wigner-Seitz radius, defined by
$\frac{4}{3}\pi a^{3}_{WS}=1/n_{0}.$ (2)
@@
The parameter $\Gamma$ is sufficient to determine the phase of a OCP in three dimensions. Plasmas with $\Gamma>1$ are called strongly coupled. The onset of fluid-like behavior, characterized by short-range oscillations in the pair correlation function, is predicted at $\Gamma\sim 2$. The oscillations in the pair correlation function grow with increasing $\Gamma$ and the system is predicted to undergo a first-order, liquid-solid phase transition to a body-centered cubic lattice at $\Gamma_{\text{crit}}\approx 172$ [252, 298, 236, 92].
@@
The classical OCP in two dimensions consists of systems of point charges that interact with the $1/r$ Coulomb repulsion but are restricted to move on a plane in a neutralizing background. The coupling constant $\Gamma$ is now determined using the Wigner-Seitz radius in two dimensions, given by $a_{\mathrm{WS,2D}}=(\pi n)^{-1/2}$ where $n$ is now the number of point charges per unit area. Here a liquid-solid phase transition at $\Gamma\approx 131(7)$ has been measured [128] in agreement with theory [116]. We note that the Mermin-Wagner theorem [210, 142], that prohibits full-fledged crystallization in two and lower dimensions, does not apply for the unscreened Coulomb repulsion, where the energy is non-additive and the interactions stabilize the ordered structure against fluctuations [45].
@@
In ion Coulomb crystals the kinetic energy can be reduced through laser cooling, thus increasing the coupling constant $\Gamma$ to achieve crystallization. Moreover, the dimensionality of the crystal is determined by controlling the relative strength of the confinement of the trap in different directions. In this way ordered structures can be realized where the dimensionality varies from one to three dimensions, including the dimensional cross-over. The Coulomb coupling $\Gamma$ for a given temperature can be estimated from Eq. (1). We take the three dimensional case and use typical trapped ion densities $10^{15}\>\text{m}^{-3}$ as reference:
$\Gamma\approx 2.69\>Z^{2}\left(\frac{n_{0}}{10^{15}\>\text{m}^{-3}}\right)^{\frac{1}{3}}\left(\frac{T}{1\,\text{K}}\right)^{-1}$ (3)
@@
where $Z$ is the charge (in elementary charges) of the trapped ion. For singly charged ions ($Z=1$), temperatures of $\sim 1\>\text{mK}$ are realized with conventional laser cooling techniques (so-called Doppler cooling), enabling Coulomb coupling parameters of several thousand, namely, an order of magnitude greater than the critical coupling parameter $\Gamma_{\text{crit}}\approx 172$ for the predicted liquid-solid phase transition. We remark that, due to the long-range nature of the Coulomb interactions, boundary effects significantly affect the resulting order. For example, observing a body-centered cubic structure in the bulk may require $N>10^{5}$ number of ions [308, 91, 133]; see Sec. II.3.1.
@@
Temperatures close to the quantum zero-point motion can be reached with more refined laser cooling techniques, see Sec. I.3. In this regime the crystals’ vibrations are quantum mechanical, but the ground state configuration is classical to a very good approximation since the typical ion densities are far from the conditions at which quantum fluctuations can melt the crystal. We provide below an estimate and refer the reader to [337] for a detailed calculation on the tunneling probability in a two-ion chain. Denoting by $\omega_{\mathrm{p}}=\left(\frac{Q^{2}n_{0}}{\epsilon_{0}m}\right)^{\frac{1}{2}}$ the ion plasma frequency, the condition for quantum melting can be cast in the form of the inequality
$\frac{\hbar}{2}\omega_{\mathrm{p}}\gtrsim\frac{1}{4\pi\epsilon_{0}}\frac{Q^{2}}{a_{\mathrm{WS}}},$ (4)
@@
where the left hand side provides an estimate of the lowest ion bound state energy due to quantum fluctuations while the right hand side is an estimate of the energy barrier impeding an ion from hopping to the neighboring site of the Wigner crystal. Equation (4) sets an upper bound on the mean interparticle distance (here expressed in terms of $a_{\mathrm{WS}}$) for the ion crystal to melt:
@@ I.2 Ion Coulomb crystals in a laboratory
It is not possible to disentangle any discussion on ion Coulomb crystals from their environment, the ion trap. Confining one-component plasmas in a laboratory requires one to either apply oscillating electric fields or to combine static electric and magnetic fields. Equilibrium, in the thermodynamic sense, can be challenging to achieve (see Sec. III.1). Below we review the basics of Paul (or rf) and Penning traps, which are the most frequently used experimental platforms. Moreover, we will briefly review optical trapping of ions by dipole traps, which has been recently realized.
@@
According to Earnshaw’s theorem [102] it is impossible to confine charged particles in vacuum exclusively by static electric fields generated from voltages applied to external electrodes. By introducing time-varying electric fields, spatially localized effective potential energy minima can, however, be realized. For atomic and not very heavy molecular species, fields in the radio frequency (rf) regime are most often applied to achieve strong confining forces as well as deep effective trap potentials (see details below). Hence, such traps are often referred to as rf traps.
Rf traps are divided into sub-categories that reflect the generic configuration of the electrodes used to apply the time-varying electrical potentials. In this section we will only discuss quadrupole-field based traps, which are mainly used in research involving Coulomb crystals. It should, however, be mentioned that crystallization in traps of higher multipole orders have been discussed in [120, 52, 48]. The standard rf trap was first realized and is named after Nobel Prize Laureate Wolfgang Paul, who invented the principle for two-dimensional confinement [244]. It is sketched in Fig. 1 and is based on three electrodes machined to have three-dimensional hyperbolic surfaces in order to create the perfect boundary conditions for an electrical quadrupole field with respect to the line of cylindrical symmetry of the electrodes ($z$-axis in Fig. 1).
[…]
With reference to the parameters defined in Fig. 1, the time-dependent electrical potential in cylindrical coordinates reads
@@
where $\rho=\sqrt{x^{2}+y^{2}}$ is the distance from the symmetry axis. From this electrical potential, Newton’s equations of motion give rise to the following classical dynamics of a single charged particle
$\frac{d^{2}u}{d\tau^{2}}+\left(a_{u}-2q_{u}\,\cos\left(2\tau\right)\right)\,u=0,\quad(u=z,\,\rho)$ (7)
@@
These equations of motion are examples of the so-called Mathieu Equations, which can lead to spatially confined motion whenever the parameters $a$ and $q$ lie within the areas of combined stability as in Fig. 2.
Figure 2: Stability diagram for standard rf (or Paul) trap. Left: Confined motion of a charged particle can only happen when its stability parameters $a_{z,\rho}$ and $q_{z,\rho}$ lie within common areas of stable motion ($z$-and $\rho$-stable). Right: The largest area of common stability (marked in red in left figure) is presented as a zoom. The parameters $\beta_{z,\rho}$ represent the ratio of the main motional frequencies of the trapped particle along the $z$ and $\rho$ directions to the rf frequency $\Omega_{rf}$. Reproduced from 122
The motion within this so-called stability diagram is generally rather complex. However, for $a,q\ll 1$ the solution to the equation of motion can very well be approximated by the trajectory
@@
with $\omega_{u}$ being given by
$\omega_{u}=\sqrt{a_{u}+\frac{q_{u}^{2}}{2}}\Omega_{\mathrm{rf}}\,.$ (12)
These solutions describe harmonic motion at frequencies $\omega_{z,\rho}$ superimposed with a much faster quiver motion at $\Omega_{rf}$. The latter motion is often referred to as micromotionmicromotion and has a small amplitude proportional to the distance of the harmonic (secular) motion from the origin (See Fig. 3).
Figure 3: Motion of a trapped particle in an rf trap. The plot shows the motion of a trapped ion along one of the principal directions ($x,y,x$) for $a_{z}=0.0$ and $q_{z}=0.3$. The red curve represents the full solution to Eq. (7), while the dark blue curve is the solution obtained by using the harmonic pseudopotential of Eq. (13). The difference between the two curves is the so-called micromotion, which has the oscillation frequency of $\Omega_{rf}$.
In many cases one can neglect the micromotion, and one is essentially left with a pure harmonic motion. To this, one can assign a potential energy often named the $\it{pseudopotential}$ which reads
@@
The sub-category of rf traps most used in experiments with ion Coulomb crystals is the linear quadrupole rf trap [255]. Ideally, it consists of four rods with hyperbolic cross-sections arranged in a parallel configuration to create a perfect two-dimensional quadrupole field for confinement in the cross-section plane (See Fig. 4 a.). Along the rod axis, confinement is obtained by applying DC voltages either to two additional electrodes placed along the central axis between the rods or to the eight end-pieces after sectioning the rods into three pieces (See Figs. 4b and c).
Figure 4: Linear trap geometries. a) Illustration of an ideal electrode geometry with hyperbolic-shaped electrodes to achieve two-dimensional trapping in the xyx-y-plane by a perfect electrical quadrupole field. This ideal geometry is often used in quadrupole mass spectrometers. b) An example of a three-dimensional linear rf trap with less ideal cylindrically shaped electrodes. Here, confinement in the xyx-y-plane is as in $a)$ by applying the same rf voltage to each of the three sections of the electrodes. Axial confinement along the $z$-axis is achieved through application of the same dc voltage $U_{end}$ to all the orange-colored electrode sections. c) Another linear trap design based on rf electrodes shaped like blades to create a smaller effective value of $r_{0}$, and hence achieve higher trap frequencies in the xyx-y-plane for a given rf voltage (see Eqs. (12) and (14)-(17). The built-in anharmonicity of this geometry will eventually limit its use to one-dimensional or smaller two- and three-dimensional ion Coulomb crystals. Reproduced from 301, 254 and 280.
The equation of motion of a charged particle in the trap geometry shown in Fig. 4b is the same as Eq. (7), but now with u=(x,y,z)u=(x,\,y,\,z) and
@@
$x$ $q_{z}=0.$ (18)
When a voltage $U_{end}$ is applied to all eight end-electrodes, $U_{dc}$ is applied to all sections of the rods in the xzx-z plane and $U_{dc}$ is applied to all sections of the rods in the yzy-z plane. $\kappa$ is a geometrical factor, which depend on the length of the central section of the rods.
In the case of $U_{dc}$=0, the stability criteria for the $x$ and $y$ motion become identical, and one can construct a stability diagram for $a=a_{x}=a_{y}=a_{z}/2$ and $q$ as presented in Fig. 5. In this case when $a$, $q$ $\ll$1, one can establish the same approximate solutions to the equation of motion as Eqs. (11) and (12) for the standard rf trap, as well as define a pseudopotential of the form shown in Eq.(13) with $\rho=\sqrt{x^{2}+y^{2}}$ and $\omega_{\rho}=\omega_{x}=\omega_{y}$.
Figure 5: Stability diagram for a linear rf trap. The grey shaded area represents the stability diagram for the specific case of $U_{DC}$=0, for which $a_{x}=a_{y}=a$ according to Eqs. (17) and (18), and $q_{x}=-q_{y}=q$. The area of the stability diagram is significantly larger than for the two-dimensional confinement of the quadrupole mass filter (the hatched areas A and B) and of the three-dimension stable trapping are of the standard rf trap (see Fig. 2). Reproduced from 85
The fact that all quadrupole traps close to the center give rise to an effective harmonic potential is a very important feature. Only such potentials will lead to a uniform particle density of three-dimensional Coulomb crystals, alike traditional crystal structures known from solid state physics. Specifically, for the two types of traps presented above, the ion number density is given by

### v2

@@ I Introduction
In a seminal article from 1934, Eugene Wigner predicted that conduction electrons in a metal at low densities can crystallize [367]. This is a peculiar situation in condensed matter, where otherwise crystallization is expected at high densities, and it results from the decay of the Coulomb interactions as the inverse of the interparticle distance $a$. For this reason, in an unscreened environment the Coulomb interactions scale with the charge density $n\sim a^{-3}$ as $E_{\rm pot}\sim n^{1/3}$. By comparison, the typical scale of the kinetic energy in a degenerate free electron gas is $E_{\rm kin}\sim\frac{h^{2}}{2ma^{2}}\sim n^{2/3}$ [190]. As a result, at low densities, the Coulomb interactions are dominant over the kinetic energy, leading to crystallization. More generally, at low densities the Wigner crystal phase is favored regardless of the particle statistics, since the overlap of single-particle wave functions is small.
@@
Wigner crystallization, here broadly defined as the formation of ordered arrays of particles that interact through Coulomb repulsion, can occur in a variety of systems. It is expected to occur in the interior of white dwarfs and in the outer crust of neutron stars [358]. It has been experimentally observed with electrons trapped on the surface of liquid helium [138], and in low-dimensional solid state devices including GaAs/GaAlAs heterojunctions [5] and carbon nanotubes [327]. The stability of Wigner crystalline islands formed in quantum dots was analyzed in [194]; Wigner molecular crystals have been reported in semiconductor moiré superlattices [374, 208]. Ordered arrays of charged, solid particles have been experimentally studied with dusty plasmas [347, 276, 217] and colloidal suspensions of polystyrene spheres in water [359, 20]. Wigner crystallization is being pursued through the formation of ultracold neutral plasmas of atomic ions immersed in a background of photo-ionized electrons [22, 203]. In many of these examples, the Coulomb repulsion of the charged objects is screened by a polarizing medium in which the objects are suspended.
[…]
@@
Wigner crystals of trapped ions, often dubbed ion Coulomb crystals, have been acquiring increasing importance for quantum technological applications [142, 27, 44, 243] thanks to advances in cooling and trapping [113, 173, 320]. These advances rely on the detailed knowledge and experimental control of the microscopic dynamics of these strongly-correlated phases of matter [173, 320]. This review reports the progress on the characterization of the equilibrium and out-of-equilibrium dynamics of ion Coulomb crystals, which has enabled this progress and has, in turn, unveiled the properties of an exotic form of condensed matter.
@@
Our starting point is the knowledge summarized in the review by Dubin and O’Neil [105] on the plasma phase and its transition to crystallization. We set our focus on the crystallized phase and discuss the insights gained over two decades of theoretical and experimental investigations. The content is complementary to recent overviews on ion Coulomb crystals [89, 348] and a wide-ranging review on strongly coupled Coulomb systems [237].
@@ I.1 Historical framework and conditions for crystallization
In 1977 Malmberg and O’Neil explained how collections of a single species of trapped charges (ions or electrons) are ideal one-component plasmas  [221]. One-component plasmas (OCP) have been the subject of detailed theoretical investigations for understanding correlations in a diverse range of physical systems [43, 18, 165, 166]. They consist of a collection of identical point charges that interact through Coulomb repulsion and are immersed in a uniform neutralizing background of given density $n_{0}$, which is assumed to be continuous and non-polarizable. In equilibrium, the point charges match their density to that of the neutralizing background. Malmberg and O’Neil showed that the forces that confine ions in a Penning trap are mathematically equivalent to a uniform, neutralizing background. More specifically, in a frame rotating with the collection of ions (or electrons) confined in a Penning trap (see Sec. I.2.2), the static thermodynamic properties of the trapped ions are the same as those for an OCP. The argument can be straightforwardly extended to ions confined in radio frequency traps in the so-called pseudo-potential approximation (see Sec. I.2.1).
@@
Here, $\epsilon_{0}$ is the permittivity of the vacuum, $Q$ is the charge and $T$ is the temperature of the ion, $k_{\mathrm{B}}$ is Boltzmann’s constant, and $a_{\mathrm{WS}}$ is the Wigner-Seitz radius, which is determined by the neutralising background density $n_{0}$ through the relation
$\frac{4}{3}\pi a^{3}_{\rm WS}=1/n_{0}\,.$ (2)
@@
The coupling constant $\Gamma$ is a measure of the ratio of the Coulomb potential energy between nearest neighbor ions (or more generally point charges) to the average kinetic energy of an ion and is sufficient to determine the phase of a OCP in three dimensions. Plasmas with $\Gamma>1$ are called strongly coupled. For a bulk (that is, infinite) OCP, the onset of fluid-like behavior, characterized by short-range oscillations in the pair correlation function, is predicted at $\Gamma\sim 2$. These oscillations become more pronounced with increasing $\Gamma$ and the system is predicted to undergo a first-order, liquid-solid phase transition to a body-centered cubic lattice at a critical $\Gamma$ [281, 333, 263, 98]. The best estimates of $\Gamma_{\text{crit}}$ lie in the range $\Gamma_{\text{crit}}\approx 173\text{--}175$ [115, 82, 105, 283].
@@
The classical OCP in two dimensions consists of systems of point charges that interact with the $1/r$ Coulomb repulsion but are restricted to move on a plane in a neutralizing background. The coupling constant $\Gamma$ is now determined using the Wigner-Seitz radius in two dimensions, given by $a_{\mathrm{WS,2D}}=(\pi n)^{-1/2}$ where $n$ is now the number of point charges per unit area. Here, a liquid-solid phase transition at $\Gamma\approx 131(7)$ has been measured [138] in agreement with theory [126]. We note that the Mermin-Wagner theorem [234, 154], that prohibits full-fledged crystallization in two and lower dimensions, does not apply for the unscreened Coulomb repulsion, where the energy is non-additive and the interactions stabilize the ordered structure against fluctuations [49].
@@
In ion Coulomb crystals the kinetic energy can be reduced through laser cooling, thus increasing the coupling constant $\Gamma$ to achieve crystallization. The dimensionality of the crystal is determined by the relative strength of the confinement of the trap in different directions. In this way ordered structures can be realized that are one-, two-, or three-dimensional. Moreover, the transition between structures of different dimensions can be experimentally studied by adjusting the relative strength of the confining potential in different directions. See Secs. II.1.4 and  II.2.4 for a detailed discussion.
The temperature required for crystallization can be estimated from Eq. (1). We take the three-dimensional case and use typical trapped ion densities of $10^{15}\>\text{m}^{-3}$ and $T=1\>\text{K}$ as a reference,
$\Gamma\approx 2.69\>Z^{2}\left(\frac{n_{0}}{10^{15}\>\text{m}^{-3}}\right)^{\frac{1}{3}}\left(\frac{T}{1\,\text{K}}\right)^{-1}\,$, (3)
@@
where $Z$ is the charge (in units of elementary charges) of the trapped ion. For singly charged ions ($Z=1$), the critical coupling parameter $\Gamma_{\text{crit}}\approx 174$ for the predicted liquid-solid phase transition for a bulk OCP is obtained with temperatures of $\sim 15\>\text{mK}$, which is more than one order of magnitude higher than the limit of conventional laser cooling techniques (so-called Doppler laser cooling, see Sec. I.3).
We emphasize that, due to the long-range nature of the Coulomb interactions, boundary effects can affect significantly the resulting spatial correlations. For example, theoretical calculations indicate that observing a body-centered cubic structure predicted for the bulk OCP may require $N>10^{5}$ ions [97, 148, 343]; see Sec. II.3.1.
@@
Temperatures close to the quantum zero-point motion can be reached with more refined laser cooling techniques, see Sec. I.3. In this regime the crystals’ vibrations are quantum mechanical, but the ground state configuration is classical to a very good approximation since the typical ion densities are far from the conditions at which quantum fluctuations can melt the crystal. Below we provide an estimate for the density required to melt an ion Coulomb crystal and note that this estimate is closely related to the problem of tunneling in a two-ion crystal chain [375]. Denoting by $\omega_{\mathrm{p}}=\left(\frac{Q^{2}n_{0}}{\epsilon_{0}m}\right)^{\frac{1}{2}}$ the ion plasma frequency, an approximate estimate of the zero-point energy of an ion localized in the crystalline phase is $\hbar\omega_{\mathrm{p}}/2$. Quantum melting is expected to occur when
$\frac{\hbar\omega_{\mathrm{p}}}{2}\gtrsim\frac{1}{4\pi\epsilon_{0}}\frac{Q^{2}}{a_{\mathrm{WS}}},$ (4)
@@
where the right side of inequality provides an estimate of the energy barrier impeding an ion from hopping to the neighboring sites of the Wigner crystal. Equation (4) sets an upper bound on the mean interparticle distance (here expressed in terms of $a_{\mathrm{WS}}$) for the ion crystal to melt:
@@ I.2 Ion Coulomb crystals in a laboratory
It is not possible to disentangle any discussion on ion Coulomb crystals from their environment, the ion trap. Confining one-component plasmas in a laboratory requires one to either apply oscillating electric fields or to combine static electric and magnetic fields. Equilibrium, in the thermodynamic sense, can be challenging to achieve (see Sec. III.1). Below we review the basics of Paul (or rf) and Penning traps, which are the most frequently used experimental platforms.
@@
According to Earnshaw’s theorem [109] it is impossible to confine charged particles in vacuum exclusively by static electric fields generated from voltages applied to external electrodes. However, by introducing time-varying electric fields, it is possible to realize spatially localized minima of the effective potential energy. For atomic and not very heavy molecular species, fields in the radio frequency (rf) regime are most often applied to achieve strong confining forces as well as deep effective trap potentials, as we detail below. Hence, such traps are often referred to as rf traps.
Figure 1: Schematics of the standard three-dimensional rf trap (Paul trap). The trap consists of a central ring electrode (named “ring” in the figure), and two so-called end cap electrodes (“Top/Bottom end-cap”). All three electrodes have hyperboloid-shaped surfaces surrounding the trap center. They produce an electrical quadrupole field when the voltage difference $\it{U_{\rm dc}+U_{\rm rf}}\cos(\Omega_{\rm rf}t)$ is applied between the ring and the end caps. The strength of the quadrupole field depends on the distances $\it{r_{0}}$ and $\it{z_{0}}$. A perfect quadrupole field is generated when $\it{r_{0}}$= $\sqrt{2}\it{z_{0}}$.
The standard rf trap is named after Nobel Prize Laureate Wolfgang Paul, who invented the principle for two-dimensional confinement [271]. It is sketched in Fig. 1 and is based on three electrodes machined to have three-dimensional hyperbolic surfaces in order to create the perfect boundary conditions for an electrical quadrupole field with respect to the line of cylindrical symmetry of the electrodes ($z$-axis in Fig. 1). With reference to the parameters defined in Fig. 1, the time-dependent electrical potential in cylindrical coordinates reads
@@
where $\rho=\sqrt{x^{2}+y^{2}}$ is the distance from the symmetry axis. Newton’s equations of motion for a single charged particle read
$\frac{d^{2}u}{d\tau^{2}}+\left(a_{u}-2q_{u}\,\cos\left(2\tau\right)\right)\,u=0,\quad(u=z,\,\rho)\,$, (7)
@@
Equations (7) are examples of the so-called Mathieu Equations. They can lead to spatially confined motion whenever the parameters $a=|a_{u}|$ and $q=|q_{u}|$ lie within the areas of combined stability as illustrated in Fig. 2. The motion within this so-called stability diagram is generally rather complex. However, for $a,q\ll 1$ the solution to the equation of motion can be well approximated by the trajectory
@@
with $\omega_{u}$ given by
$\omega_{u}=\frac{1}{2}\sqrt{a_{u}+\frac{q_{u}^{2}}{2}}\Omega_{\mathrm{rf}}\,.$ (12)
These solutions describe harmonic motion at frequencies $\omega_{z,\rho}$ superimposed with a much faster quiver motion at $\Omega_{\mathrm{rf}}$. The quiver motion is often referred to as micromotionmicromotion and has a small amplitude proportional to the distance of the harmonic (secular) motion from the origin (See Fig. 3). In many experimental situations, one can neglect the micromotion and is essentially left with pure harmonic motion. In such situations, the term proportional to $q_{u}$ in Eq.(11) can be discarded and one assigns a potential energy to the trap, often referred to as the $\it{pseudopotential}$, which reads
@@
Figure 2: Stability diagram for standard rf trap. Left: Confined motion of a charged particle can only happen when its stability parameters $a_{z,\rho}$ and $q_{z,\rho}$ lie within common areas of stable motion ($z$- and $\rho$-stable). Right: Zoom on the largest area of common stability (marked in red in left panel). The parameters $\beta_{j}$ represent the ratio between the main motional frequencies of the trapped particle along the $j$ direction ($j=z,\rho)$ and the rf frequency $\Omega_{\rm rf}/2$. Reproduced from 132
Figure 3: (Color online) Motion of a trapped particle in an rf trap. The plot shows the trajectory of a trapped ion along one of the principal directions ($x,y,x$) for $a_{z}=0.0$ and $q_{z}=0.3$. The oscillatory red (solid) curve represents the full solution to Eq. (7), while the dark blue (dashed) curve is the solution obtained by using the harmonic pseudopotential of Eq. (13). The difference between the two curves is the so-called micromotion at the oscillation frequency $\Omega_{\rm rf}$.
The types of rf traps most used in experiments with ion Coulomb crystals are linear quadrupole rf traps [285]. Ideally, a linear quadrupole rf trap consists of four rods with hyperbolic cross-sections arranged in a parallel configuration to create a perfect two-dimensional quadrupole field for confinement in the cross-section plane (See Fig. 4 a.). Along the rod axis, confinement is obtained by applying DC voltages either to two additional electrodes placed along the central axis between the rods or to the eight end-pieces after sectioning the rods into three pieces (See Figs. 4b and c).
Figure 4: Linear trap geometries. a) Illustration of an ideal electrode geometry with hyperbolic-shaped electrodes to achieve two-dimensional trapping in the xyx-y plane by a perfect electrical quadrupole field. This ideal geometry is often used in quadrupole mass spectrometers. b) An example of a three-dimensional linear rf trap with cylindrically-shaped electrodes. Here, one realizes a confinement in the xyx-y plane like in $a)$ by applying the same rf voltage to each of the three sections of the electrodes. Axial confinement along the $z$-axis is achieved through application of the same dc voltage $U_{\rm end}$ to all the orange-colored electrode sections. c) Linear trap based on rf electrodes shaped like blades. This design creates a smaller effective value of $r_{0}$, and hence achieves higher trap frequencies in the xyx-y plane at a given rf voltage (see Eqs. (12) and (14)-(17). The built-in anharmonicity of this geometry eventually limits its use to one-dimensional ion Coulomb crystals or to small structures of two or three dimensions. Reproduced from 336, 284 and 314.
For the trap geometry of Fig. 4, the equations of motion of a charged particle have the same form as Eq. (7), but now with u=(x,y,z)u=(x,\,y,\,z) and
@@
$q_{z}=0.$ (18)
When a voltage $U_{\rm end}$ is applied to all eight end-electrodes, $U_{\mathrm{\rm dc}}$ is applied to all sections of the rods in the xzx-z plane and $U_{\mathrm{\rm dc}}$ is applied to all sections of the rods in the yzy-z plane. $\kappa$ is a geometrical factor, which depends on the length of the central section of the rods.
For $U_{\mathrm{\rm dc}}$=0, the stability criteria for the $x$ and $y$ motion become identical, and one can construct a stability diagram for $a=a_{x}=a_{y}=a_{z}/2$ and $q$ as presented in Fig. 5. In this case when $a$, $q$ $\ll$1, one can establish the same approximate solutions to the equation of motion as Eqs. (11) and (12) for the standard rf trap, as well as define a pseudopotential of the form shown in Eq.(13) with $\rho=\sqrt{x^{2}+y^{2}}$ and $\omega_{\rho}=\omega_{x}=\omega_{y}$.
Figure 5: Stability diagram for a linear rf trap. The grey shaded area represents the stability region for the specific case of $U_{\mathrm{dc}}$=0, for which $a_{x}=a_{y}=a$ according to Eqs. (17) and (18), and $q_{x}=-q_{y}=q$. The area of the stability diagram is significantly larger than for the two-dimensional confinement of the quadrupole mass filter (the hatched areas A and B) and of the three-dimension stable trapping are of the standard rf trap (see Fig. 2). Reproduced from 91
The fact that all quadrupole traps close to the center give rise to an effective harmonic potential is a very important feature. Only such potentials will lead to a uniform particle density of three-dimensional Coulomb crystals, like traditional crystal structures known from solid state physics. In this case, for the linear rf trap presented above, the ion number density is given by
@@
For trap parameters $U_{\mathrm{rf}}\sim 100-1000\,\mathrm{V}$, $r_{0}\sim 1-5\,\mathrm{mm}$, and $\Omega_{\mathrm{rf}}\sim 2\pi\times 10\,\mathrm{MHz}$, one typically reaches ion densities $n_{0}\sim 10^{14}-10^{15}\,\mathrm{m}^{-3}$.